OpenStudy (anonymous):

Please help me answer the last part of this question below, the explain if the two graphs are the same, THANKS so much!! show that the solution of: dy/dx = -((1+y^2)/(1+x^2)) y(0) = -1 is the inverse tan (x) + inverse tangent (y) = -pi/4 AND show that the above can be written : y= ((x+1)/(x-1)) And are the graphs of: inverse tan(x) + inverse tan (y) = -pi/4 AND y=((x+1)/(x-1)) from above the same? Explain.

OpenStudy (dumbcow):

use identity of angle addition with tangent \[\tan (\tan^{-1} x + \tan^{-1} y) = \tan(-\pi/4)\]

OpenStudy (dumbcow):

\[\tan (a+b) = \frac{\tan a + \tan b}{1 - \tan a \tan b}\] \[\rightarrow \frac{x+y}{1-xy} = -1\]

OpenStudy (dumbcow):

solving for y will give \[y = \frac{x+1}{x-1}\] and yes i believe both graphs would be the same

OpenStudy (anonymous):

Do you know why they are the same though? I dont know how to explain it